# Impact of heterogeneity on infection probability: Insights from single-hit dose-response models

### v1

@@ Abstract
The process of infection of a host is complex, influenced by factors such as microbial variation within and between hosts as well as differences in dose across hosts. This study uses dose-response and microbial growth models to delve into the impact of these factors on infection probability. It is rigorously demonstrated that within-host heterogeneity in microbial infectivity enhances the probability of infection. The effect of infectivity and dose variation between hosts is studied in terms of the expected value of the probability of infection. General analytical findings, derived under the assumption of small infectivity, reveal that both types of heterogeneity reduce the expected infection probability. Interestingly, this trend appears consistent across specific dose-response models, suggesting a limited role for the small infectivity condition. Additionally, the vital dynamics behind heterogeneous infectivity are investigated with a microbial growth model which enhances the biological significance of single-hit dose-response models. Testing these mathematical predictions inspire new and challenging laboratory experiments that could deepen our understanding of infections.
@@
Probability of infection , Heterogeneity , Dose-response models , Single-Hit Model , Microbial growth dynamics
@@
[inst1]organization=Institute for Complex Systems and Mathematical Biology, School For Natural and Computing Sciences, SUPA,addressline=University of Aberdeen, city=Aberdeen, postcode=AB24 3UE, country=United Kingdom
@@ 1 Introduction
Estimating the probability of infection following exposure to a pathogenic microbe is a crucial step in quantitative microbial risk assessment [8]. Heterogeneity associated with the host-microbe interaction within hosts and differences between hosts in terms of the ingested dose and susceptibility, are expected to influence the chances of infection. However, a systematic understanding of the role of heterogeneity on infection is lacking. Interpreting infection as a biological invasion could guide intuition on the role of heterogeneity in the probability of infection. Different types of heterogeneity have been considered by ecological and epidemiological models for the probability of biological invasions [9, 10, 11, 12, 13]. Such models, however, suggest that depending on the type of heterogeneity, the probability of invasion could either be enhanced of reduced.
@@
Ingestion of a single pathogenic microbe is sufficient to induce infection in a host (i.e., no combined action of two or more microbes is required for infection, provided proliferation from a single microbe occurs).
@@
If a host ingests a dose containing several pathogenic microbes, the survival of any microbe within the host is independent of the survival of any other microbe.
@@
The systematic investigation of whether each type of heterogeneity enhances or diminishes the probability of infection has not been systematically undertaken in previous studies. Furthermore, a link between heterogeneous infectivity and microbial growth mechanisms (e.g., microbial rates of birth and death) has yet to be established. Advancing in these two directions constitutes the central motivation of the present work.
@@
Concerning the second direction mentioned above, the growth dynamics of microbes have been described with birth-death models that have been related to dose-response models [28, 29]. However, heterogeneous infectivity was not incorporated into these models [30], and for instance, a mechanistic interpretation of the parameters $\alpha$ and $\beta$ of the beta-Poisson model has not been provided.
@@ 2 General single-hit dose-response model
Let as assume a population of $H$ hosts such that each host ingest (or is in contact with) $n_{h}$ pathogenic microbes. The infectivity of a microbe $i$ is quantified by the probability $x_{i,h}$ that the microbe grows and infects the host. From assumption 1 in the introduction, any microbe with $x_{i,h}>0$ can cause infection. Following assumption 2, the probability that a host is infected is
@@
$P_{h}\left(\{x_{i}\}_{i=1}^{n},n\right)=1-(1-\bar{x})^{n}\exp{\left(-n\sum_{k=2}^{\infty}\frac{M_{k}}{k(1-\bar{x})^{k}}\right)}~{},$ (3)
@@ 4.1 General results for the expected probability of infection
(b) The between-host variation of infectivity is described by a random variable with expectation $\mu_{x}\in[0,1]$ and variance $v_{x}\in[0,1/4]$.
(c) The dose ingested by different hosts is a random variable with expectation $\mu_{n}\geq 1$ (i.e. at least one microbe is ingested) and variance $v_{n}\geq 0$.
@@
It is now convenient to express the expected value in Eq. (13) as ${\text{E}}_{x}({\text{E}}_{n}((1-x)^{n}))=\exp(\ln(1+y))$ and use the expansion $\ln{(1+y)}=y-y^{2}/2+O(y^{3})$ which, after neglecting $O(x^{3})$ terms, leads to the following approximation:
@@
According to Eq. (15), $\hat{P}$ decreases with increasing $v_{n}$ since $\mu_{x}^{2}>0$. $\hat{P}$ also decreases with increasing $v_{x}$ since $\mu_{n}^{2}-\mu_{n}+v_{n}\geq 0$ for any reasonable dose with $\mu_{n}\geq 1$ and $v_{n}\geq 0$. ∎
@@
In the limit of small infectivity, the slope of $\hat{P}$ evaluated at the median dose is an increasing function of $\mu_{x}$ and a decreasing function of $v_{x}$ for any dose distribution.
@@
$\left.\frac{\text{d}\hat{P}}{\text{d}\mu_{n}}\right|_{\mu_{n}=\mu_{50}}=\frac{\mu_{x}}{2}\left(1-\frac{\mu_{x}}{2}\right)-\frac{v_{x}}{2}\left(\mu_{50}-\frac{1}{2}\right)~{}.$ (16)
This equation trivially proves the corollary for any reasonable parameters $\mu_{50}>1/2$ and $\mu_{x}\leq 1$. ∎
@@ 4.2 Models of type $A^{\prime}$: Homogeneous infectivity and heterogenous dose
Models of type $A^{\prime}$ extend Model A to include heterogeneity of type III (see Figure 1). Specifically, it is posited that the doses ingested by different hosts are independent and identically distributed (iid) random variables drawn from a probability distribution $p_{n}^{\text{d}}(\bm{\xi})$, where $\bm{\xi}$ are the distribution parameters. The expected probability of infection can then be expressed in terms of $P^{(A)}(x,n)$ as follows:
@@ 4.2.2 Example $A^{\prime}_{2}$: Negative binomial dose distribution
The Poisson distribution assumes $\mu_{n}=v_{n}$ and does not allow the effect of the effect of $v_{n}$ on the expected probability of infection to be independently analysed from $\mu_{n}$. In fact, doses can be characterised by a larger dispersion than that allowed by a Poisson distribution (i.e. $v_{n}>\mu_{n}$[34, 35, 21]. Over-dispersion can be modelled by assuming a negative binomial distribution, $n_{h}\sim\text{NB}(r,p)$, with parameters $r>0$ and $p\in[0,1]$. The corresponding probability function is
@@
which results in the following expected probability of infection:
@@
Note that the condition $r>0$ implies $v_{n}\geq\mu_{n}$ for the negative binomial distribution. The smallest possible dispersion for a negative binomial distribution corresponds to $v_{n}=\mu_{n}$ which is the condition defining a Poisson distribution. The following proposition shows that the expected probability of invasion in examples $A^{\prime}_{1}$ and $A^{\prime}_{2}$ are identical when $v_{n}=\mu_{n}$.
@@
$\lim_{t\rightarrow\infty}P^{(A^{\prime}_{2})}(x,\mu_{n},r)=1-\lim_{t\rightarrow\infty}\left(1+\frac{-x\mu_{n}}{-r}\right)^{-r}=1-e^{-x\mu_{n}}~{}.$ (23)
@@
$f(y,r)$ has only one zero and is located at $y=0$ for any $r>0$. To show this, it is convenient to define a new variable $z=\frac{r+y}{er}$ such that $f(y,r)=0$ reduces to $z\ln z=e^{-1}$. The solution to this equation is $z=-e^{-1}$ which corresponds to $y=0$.
@@
Following these results for $f(y,r)$ must be non-positive for any $y>0$ and $r>0$. ∎
@@
For a fixed mean dose, the expected probability of invasion of model $A^{\prime}_{2}$ is a decreasing function of the dose variance.
@@ 4.4.1 Example C1: Beta-distributed between-host infectivity
Note that the conditions $\alpha,\beta>0$ impose an upper bound for the variance $v_{x}<v_{\max}(\mu_{x})\equiv\mu_{x}(1-\mu_{x})$.
@@
This regime corresponds to $\alpha,\beta<1$. The infectivity PDF $\rho_{h}^{(C_{1})}$ diverges for $x=0$ and $x=1$ (see Fig. 3(a)). This corresponds to populations of hosts that tend to have a bimodal distribution of infectivities: The microbe can be highly infective for some hosts whereas it tends to be harmless for others. The relative proportion of hosts at risk of infection is larger when $\beta<\alpha$ as this makes the divergence of $\rho_{h}^{(C_{1})}$ at $x=1$ dominant over the one at $x=0$.
@@
Figure 3: Probability density function of the infectivity for model $C_{1}$ (red dashed line, Eq. (31)) and the population growth model proposed in Sec. 5.0.2 (black line, Eq. (56)). Different panels illustrate the dependence of $\rho_{h}^{(C_{1})}(x;\alpha,\beta)$ on infectivity in each of the regimes described in the text: (a) Regime (i) with $(\alpha,\beta)=(0.5,0.25)$, (b) regime (ii) with $(\alpha,\beta)=(1.2,0.36)$, (c) regime (iii) with $(\alpha,\beta)=(0.89,1.32)$, and (d) regime (iv) with $(\alpha,\beta)=(1.5,1.8)$. When parametrised by the mean and variance of the effective microbe mortality considered in the population growth model of Sec. 5.0.2, the four panels correspond to: (a) $(\mu_{\lambda},v_{\lambda})=(0.5,1.0)$, (b) $(\mu_{\lambda},v_{\lambda})=(0.3,0.25)$, (c) $(\mu_{\lambda},v_{\lambda})=(1.5,1.7)$ and (d) $(\mu_{\lambda},v_{\lambda})=(1.2,0.8)$.
Figure 4: Location of the four regimes for the shape of the PDF $\rho_{h}^{(C_{1})}(x;\alpha,\beta)$ in the space spanned by the mean and variance of the infectivity.
A fully analytical study of the trends of the probability of infection $P^{(C_{1})}$ when varying $\mu_{x}$ and $v_{x}$ is challenging and a graphical analysis will be presented instead. Fig. 5(a) shows the dependence of $P^{(C_{1})}$ on the mean infectivity for two values of the variance (see red and black continuous lines). For given $v_{x}$, the condition $v_{x}<\mu_{x}(1-\mu_{x})$ implies that $P^{(C_{1})}$ is only defined for $\mu_{x}\in(\mu_{x}^{-},\mu_{x}^{+})$, where $\mu_{x}^{\pm}(v_{x})=(1\pm\sqrt{1-4v_{x}})/2$. At the boundaries $\mu_{x}^{\pm}$, the probability of infection satisfies
@@
Figure 5: Expected probability of infection for examples $C_{1}$ with dose $n=5$ (continuous lines, Eq. (32)) and $\tilde{C_{1}}$ (dashed lines, Eq. (38)). Panel (a) shows the dependence of $P^{(C_{1})}$ and $P^{(\tilde{C_{1}})}$ on the mean infectivity, $\mu_{x}$, for two values of the variance, $v_{x}$, as marked by the legend. The dotted diagonal line indicates the bounds and (b) $v_{x}$
@@
Comparing the continuous and dashed lines in Fig. (5)(b) reveals that $P^{(\tilde{C_{1}})}$ offers a better approximation to $P^{(C_{1})}$ for small values of $v_{x}$ and $\mu_{x}$ (compare the good agreement between the red curves for small $v_{x}$ with the poorer agreement between the black curves for small $v_{x}$). In the limit of vanishing variance, the approximate probability approaches the value
@@ 4.5.1 Example $C_{1}^{\prime}$: Beta-distributed infectivity and Poisson-distributed dose
Here, ${}_{1}F_{1}$ is the Kummer confluent function [36]. This is the exact beta-Poisson model proposed by Furumoto and Mickey [16]. In the same work, the authors derived the beta-Poisson approximation,

### v2

@@ Abstract
The process of infection of a host is complex, influenced by factors such as microbial variation within and between hosts as well as differences in dose across hosts. This study uses dose-response and within-host microbial infection models to delve into the impact of these factors on infection probability. It is rigorously demonstrated that within-host heterogeneity in microbial infectivity enhances the probability of infection. The effect of infectivity and dose variation between hosts is studied in terms of the expected value of the probability of infection. General analytical findings, derived under the assumption of small infectivity, reveal that both types of heterogeneity reduce the expected infection probability. Interestingly, this trend appears consistent across specific dose-response models, suggesting a limited role for the small infectivity condition. Additionally, the vital dynamics behind heterogeneous infectivity are investigated with a within-host microbial growth model which enhances the biological significance of single-hit dose-response models. Testing these mathematical predictions inspire new and challenging laboratory experiments that could deepen our understanding of infections.
@@
Probability of infection , Heterogeneity , Dose-response models , Single-Hit Model , Within-host microbial infection dynamics
journal: Mathematical Biosciences
@@
[inst1]organization=School For Natural and Computing Sciences, SUPA,addressline=University of Aberdeen, city=Aberdeen, postcode=AB24 3UE, country=United Kingdom
@@ 1 Introduction
Estimating the probability of infection following exposure to a pathogenic microbe is a crucial step in quantitative microbial risk assessment [8]. Heterogeneity associated with the host-microbe interaction within hosts and differences between hosts in terms of the ingested dose and susceptibility, are expected to influence the chances of infection. However, a systematic understanding of the role of heterogeneity on infection is lacking. Interpreting infection as a biological invasion could guide intuition on the role of heterogeneity in the probability of infection. Different types of heterogeneity have been considered by ecological and epidemiological models for the probability of biological invasions [9, 10, 11, 12, 13]. Such models, however, suggest that depending on the type of heterogeneity, the probability of infection could either be enhanced of reduced.
@@
Ingestion of a single pathogenic microbe is sufficient to infect a host.
@@
If a host ingests a dose containing multiple pathogenic microbes, the event of one microbe causing infection is independent of the event of another microbe causing infection.
@@
The systematic investigation of whether each type of heterogeneity enhances or diminishes the probability of infection has not been systematically undertaken in previous studies. Furthermore, a link between heterogeneous infectivity and microbial infection mechanisms (e.g., microbial rates of birth and death) has yet to be established [28]. Advancing in these two directions constitutes the central motivation of the present work.
@@
Concerning the second direction mentioned above, the growth dynamics of microbes within a host have been described with birth-death models that have been related to dose-response models [29, 30]. However, heterogeneous infectivity was not incorporated into these models [28], and for instance, a mechanistic interpretation of the parameters $\alpha$ and $\beta$ of the beta-Poisson model has not been provided.
@@ 2 General single-hit dose-response model
Let us assume a population of $H$ hosts such that each host ingest (or is in contact with) $n_{h}$ pathogenic microbes. The infectivity of a microbe $i$ is quantified by the probability $x_{i,h}$ that the microbe infects the host. From assumption 1 in the introduction, any microbe with $x_{i,h}>0$ can cause infection. Following assumption 2, the probability that a host is infected is
@@
$P_{h}\left(\{x_{i}\}_{i=1}^{n},n\right)=1-(1-\bar{x})^{n}\exp{\left(-n\sum_{k=1}^{\infty}\frac{M_{k}}{k(1-\bar{x})^{k}}\right)}~{},$ (3)
@@ 4.1 General results for the expected probability of infection
(b) The between-host variation of infectivity is described by a continuous random variable defined on $[0,1]$ with expectation $\mu_{x}$ and variance $v_{x}$.
(c) The dose ingested by different hosts is a discrete random variable taking values in $\mathbb{Z}^{+}$ (i.e., at least one microbe is ingested), with expectation $\mu_{n}$ and variance $v_{n}$.
@@
It is now convenient to express the expected value in Eq. (13) as ${\text{E}}_{x}({\text{E}}_{n}((1-x)^{n}))=\exp(\ln(1+y))$ and use the expansion $\ln{(1+y)}=y-y^{2}/2+O(y^{3})$. After introducing the expression (14) into this expansion and neglecting $O(x^{3})$ terms in the resulting function of $x$, one obtains the following approximation:
@@
According to Eq. (15), $\hat{P}$ decreases with increasing $v_{n}$ for any $\mu_{x}>0$. $\hat{P}$ also decreases with increasing $v_{x}$ since $\mu_{n}^{2}-\mu_{n}+v_{n}\geq 0$ for any reasonable dose with $\mu_{n}\geq 1$ and $v_{n}\geq 0$. ∎
@@
In the limit of small infectivity, the slope of $\hat{P}$ evaluated at the median dose—defined as the dose $\mu_{n}=\mu_{50}$ for which the expected probability of infection is 50%—is an increasing function of $\mu_{x}$ and a decreasing function of $v_{x}$ for any dose distribution.
@@
$\left.\frac{\text{d}\hat{P}}{\text{d}\mu_{n}}\right|_{\mu_{n}=\mu_{50}}=\frac{\mu_{x}}{2}\left(1+\frac{\mu_{x}}{2}\right)-\frac{v_{x}}{2}\left(\mu_{50}-\frac{1}{2}\right)~{}.$ (16)
This equation trivially proves the corollary for any reasonable median dose $\mu_{50}>1/2$. ∎
@@ 4.2 Models of type $A^{\prime}$: Homogeneous infectivity and heterogenous dose
Models of type $A^{\prime}$ extend Model A to include heterogeneity of type III (see Figure 1). Specifically, it is posited that the doses ingested by different hosts are independent and identically distributed (iid) random variables drawn from a probability mass function $p_{n}^{\text{d}}(\bm{\xi})$, where $\bm{\xi}$ are the distribution parameters. The expected probability of infection can then be expressed in terms of $P^{(A)}(x,n)$ as follows:
@@ 4.2.2 Example $A^{\prime}_{2}$: Negative binomial dose distribution
The Poisson distribution assumes $\mu_{n}=v_{n}$ and does not allow the effect of $v_{n}$ on the expected probability of infection to be independently analysed from $\mu_{n}$. In fact, doses can be characterised by a larger dispersion than that allowed by a Poisson distribution (i.e. $v_{n}>\mu_{n}$[34, 35, 21]. Over-dispersion can be modelled by assuming a negative binomial distribution, $n_{h}\sim\text{NB}(r,p)$, with parameters $r>0$ and $p\in[0,1]$. The corresponding probability function is
@@
which, after introducing into (18), results in the following expected probability of infection:
@@
Note that the condition $r>0$ implies $v_{n}\geq\mu_{n}$ for the negative binomial distribution. The smallest possible dispersion for a negative binomial distribution corresponds to $v_{n}=\mu_{n}$ which is the condition defining a Poisson distribution. The following proposition shows that the expected probability of infection in examples $A^{\prime}_{1}$ and $A^{\prime}_{2}$ are identical when $v_{n}=\mu_{n}$.
@@
$\lim_{r\rightarrow\infty}P^{(A^{\prime}_{2})}(x,\mu_{n},r)=1-\lim_{r\rightarrow\infty}\left(1+\frac{-x\mu_{n}}{-r}\right)^{-r}=1-e^{-x\mu_{n}}~{}.$ (23)
@@
$f(y,r)$ has only one zero and is located at $y=0$ for any $r>0$. To show this, it is convenient to define a new variable $z=\frac{r+y}{er}$ such that $f(y,r)=0$ reduces to $z\ln z=-e^{-1}$. The solution to this equation is $z=e^{-1}$ which corresponds to $y=0$.
@@
Following these results, $f(y,r)$ must be non-positive for any $y>0$ and $r>0$. ∎
@@
For a fixed mean dose, the expected probability of infection of model $A^{\prime}_{2}$ is a decreasing function of the dose variance.
@@ 4.4.1 Example C1: Beta-distributed between-host infectivity
Note that the conditions $\alpha,\beta>0$ impose an upper bound for the variance:
$v_{x}<v_{\max}(\mu_{x})\equiv\mu_{x}(1-\mu_{x})~{}.$ (34)
In particular, this implies an absolute upper bound for the variance: $v_{x}<1/4$.
@@
This regime corresponds to $\alpha,\beta<1$. The infectivity PDF $\rho_{h}^{(C_{1})}$ diverges for $x=0$ and $x=1$ (see Fig. 3(a)). This describes populations of hosts that tend to have a bimodal distribution of infectivities: The microbe can be highly infective for some hosts whereas it tends to be harmless for others. The relative proportion of hosts at risk of infection is larger when $\beta<\alpha$ as this makes the divergence of $\rho_{h}^{(C_{1})}$ at $x=1$ dominant over the one at $x=0$.
@@
Figure 3: Probability density function of the infectivity for model $C_{1}$ (dashed line, Eq. (31)) and the within-host microbial infection model proposed in Sec. 5.0.2 (continous line, Eq. (58)). Different panels illustrate the dependence of $\rho_{h}^{(C_{1})}(x;\alpha,\beta)$ on infectivity in each of the regimes described in the text: (a) Regime (i) with $(\alpha,\beta)=(0.5,0.25)$, (b) regime (ii) with $(\alpha,\beta)=(1.2,0.36)$, (c) regime (iii) with $(\alpha,\beta)=(0.89,1.32)$, and (d) regime (iv) with $(\alpha,\beta)=(1.5,1.8)$. When parametrised by the mean and variance of the effective microbe mortality considered in the microbial infection model of Sec. 5.0.2, the four panels correspond to: (a) $(\mu_{\lambda},v_{\lambda})=(0.5,1.0)$, (b) $(\mu_{\lambda},v_{\lambda})=(0.3,0.25)$, (c) $(\mu_{\lambda},v_{\lambda})=(1.5,1.7)$ and (d) $(\mu_{\lambda},v_{\lambda})=(1.2,0.8)$.
Figure 4: Location of the four regimes for the shape of the infectivity PDF $\rho_{h}^{(C_{1})}(x;\alpha,\beta)$ in the space spanned by the mean and variance of the infectivity.
A fully analytical study of the trends of the probability of infection $P^{(C_{1})}$ when varying $\mu_{x}$ and $v_{x}$ is challenging and a graphical analysis will be presented instead. The continuous lines in Fig. 5(a) show the dependence of $P^{(C_{1})}$ on the mean infectivity for two values of the variance. For given $v_{x}$, the condition (34) implies that $P^{(C_{1})}$ is only defined for $\mu_{x}\in(\mu_{x}^{-},\mu_{x}^{+})$, where
$\mu_{x}^{\pm}(v_{x})=(1\pm\sqrt{1-4v_{x}})/2~{}.$ (36)
At the boundaries $\mu_{x}^{\pm}$, the probability of infection satisfies
@@
Figure 5: Expected probability of infection for examples $C_{1}$ with dose $n=5$ (continuous lines, Eq. (32)) and $\tilde{C_{1}}$ (dashed lines, Eq. (40)). Panel (a) shows the dependence of $P^{(C_{1})}$ and $P^{(\tilde{C_{1}})}$ on the mean infectivity, $\mu_{x}$, for two values of the variance, $v_{x}$, as marked by the legend. The dotted diagonal line indicates the bounds for $P^{(C_{1})}$ given by Eq. (37). Panel (b) shows the dependence of $P^{(C_{1})}$ and $P^{(\tilde{C_{1}})}$ on the infectivity variance $v_{x}$. The dotted line shows the limit of $P^{(C_{1})}$ as $v_{x}\nearrow v_{\max}(\mu_{x})$ (cf. Eqs. (34) and (39)).
@@
Comparing the continuous and dashed lines in Fig. (5)(b) reveals that $P^{(\tilde{C_{1}})}$ offers a better approximation to $P^{(C_{1})}$ for small values of $v_{x}$ and $\mu_{x}$ (compare the good agreement between the curves for small $v_{x}$ with the poorer agreement between the black curves for small $v_{x}$). In the limit of vanishing variance, the approximate probability approaches the value
@@ 4.5.1 Example $C_{1}^{\prime}$: Beta-distributed infectivity and Poisson-distributed dose
Here, ${}_{1}F_{1}$ is the Kummer confluent function [36]. Eq. (48) is the exact beta-Poisson model proposed by Furumoto and Mickey [16]. In the same work, the authors derived the beta-Poisson approximation,